OpenStudy (openstudyuser):

Linear Programming: Carpentry: Emilio has a small carpentry shop where he makes large and small bookcases. His profit on a large bookcase is $80 and on a small book case it is $50. It takes Emilio 6 hours to make a large bookcase and 2 hours to make a small one. He can spend only 24 hours each week on his carpentry work. He must make at least two of each size each week. What is the maximum weekly profit?

OpenStudy (anonymous):

Wait, so we don't know the profit of a small book case?

OpenStudy (openstudyuser):

Hold up, i screwed the question, ima edit it now

OpenStudy (openstudyuser):

There.

OpenStudy (anonymous):

So let's write out some equations. \[p=80l+50s\]\[6l+2s \leq 24\]\[l\geq 2\]\[ w\geq2\]

OpenStudy (anonymous):

Why did you put 'Linear Programming' up there?

OpenStudy (openstudyuser):

Because that is what this is. And i dont understand how you got the equations.

OpenStudy (anonymous):

\(p\) is just our profit equation

OpenStudy (anonymous):

Oops, \(w\geq 2\) should be \(s \geq 2\)

OpenStudy (openstudyuser):

lol, i was like where did the w come out of.

OpenStudy (anonymous):

Where as \(s\) and \(l\) are just the number of each bookcase.

OpenStudy (openstudyuser):

Could we use x and y if you dont mind, i can understand better with those variables.

OpenStudy (anonymous):

Is this a programming course?

OpenStudy (openstudyuser):

Its precal

OpenStudy (anonymous):

Do you know how to graph inequalities?

OpenStudy (openstudyuser):

Yes. We were given a graph to plot points on too.

OpenStudy (anonymous):

If you let \(l=x\) and \(s=y\) then you can graph and get all the possible solution sets which fit the inequalities... but that still won't give you the max.

OpenStudy (anonymous):

Tough question for precalc

OpenStudy (openstudyuser):

lol, so when i graph these, what do i do?

OpenStudy (anonymous):

What I would do actually is.... let \(x = l-2\) and \(y=s-2\). Then we get: \[6(x+2) + 2(y+2) \leq 24\]\[ x\geq 0\]\[y \geq 0\]

OpenStudy (openstudyuser):

How did you do that?

OpenStudy (anonymous):

Then we get: \[6x + 12 +2y + 4 \leq 24 \Rightarrow 6x+2y \leq 8\]

OpenStudy (anonymous):

Well if I set \(x = l-2\) then \(l=x+2\) if we want to replace \(l\) with \(x\)

OpenStudy (openstudyuser):

Wait, in still a little confused, how did you have x=1-2 and y=s-2?

OpenStudy (anonymous):

I'm sorry for confusing you... Since \(x\) and \(y\) are variables we are bringing into this equation, we can set them equal to whatever we like.

OpenStudy (anonymous):

The reason I chose \(l-2\) is because \(l \leq 2 \) gives us \(l-2\leq 0 \)

OpenStudy (anonymous):

And \(\leq 0\) is easier to deal with than \(\leq 2 \)

OpenStudy (openstudyuser):

Cant we just graph those?

OpenStudy (anonymous):

Sure

OpenStudy (openstudyuser):

Its much easier to graph than solve it all.

OpenStudy (anonymous):

Depends on the person, but do whatever you think is easier.

OpenStudy (openstudyuser):

\[x \le2\] is pretty easy to graph. So i would graph that alongside the other equations?

OpenStudy (anonymous):

Basically, graph all of the inequalities.

OpenStudy (openstudyuser):

\[6x+2y \le24\] How did you get that?

OpenStudy (anonymous):

Solve for \(y\).

OpenStudy (openstudyuser):

No, i mean, how did you recieve that equation?

OpenStudy (anonymous):

Because the large ones take 6 hours, the small ones take 2 hours and you can only spend 24 hours a week.

OpenStudy (openstudyuser):

Ohhh okayy, got it. Thanks for the explanation. So once we graph this, it should make a shape, in which we will take the vertices, and plug them into the Profit equation?

OpenStudy (anonymous):

You will get a shape, you want to take every coordinate in the shape (that is a whole number) and plug them into the profit equation.

OpenStudy (openstudyuser):

And once i get that coordinate, i will have to plug each one in to find the largest amount which will be the maximum weekly profit right?

OpenStudy (anonymous):

Yes

OpenStudy (openstudyuser):

Thanks soo much.